OpenStudy (anonymous):

What is the solution set for |–2x + 16| – 9 = –3

OpenStudy (anonymous):

hi there

OpenStudy (anonymous):

heyy

OpenStudy (anonymous):

5 and 11

OpenStudy (anonymous):

\[|-2x + 16 | = 6\]

OpenStudy (anonymous):

case 1: -2x + 16 = 6 gives x = 5

OpenStudy (anonymous):

case 2: -2x + 16 = -6 gives x = 11

OpenStudy (cwrw238):

|–2x + 16| – 9 = –3 |–2x + 16| = 6 as the expression in modulus sign can be positive or negative we have: -2x + 16 = 6 2x = 10 x = 5 OR -2x + 16 = -6 2x = 22 x=11

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

there u hv a detailed answer vy cwr :D